If one pump takes 12 minutes to fill a tank and another pump takes 20 minutes to fill the same size tank, how long would it take both pumps working together to fill one of the tanks?
Answer by feliz1965(151)

12 minutes is 1/5 of an hour.
20 minutes is 1/3 of an hour.
working together = 1/x
1/5 + 1/3 = 1/x
Our LCD = 15x
Multiply each term by the LCD to remove the fractions.
3x + 5x = 15
8x = 15
x = 15/8
x = 1_(7/8)